# How Far are VLMs from Visual Spatial Intelligence? A Benchmark-Driven Perspective

### v1

@@ 4.3.1 Environment Planning
Shortcut Discovery. After observing a video walkthrough from the start to the goal (with unnecessary steps), this task requires VLMs to plan a never-seen shortcut from the start to the goal [199]. The challenges of this task lie in (1) overall understanding of the environment, (2) interactive planning of actions given previous observations.
Route Retracing. In SPACE [199], this task requires VLMs to retrace the route from the start after observing a video from the start to the goal. The route in the video is always the shortest path to the goal. Similar to shortcut discovery, the challenge of this task also lies in the interactive planning of actions.
Map Sketching. After observing a video walkthrough from the start to the goal, map sketching [199] requires VLMs to pick the best map representation of the environment from multiple options. The map contains landmarks, including the start and goal, with different spatial relationships. The challenges of this task lie in (1) recognition of critical landmarks, (2) comprehension of spatial relationships among landmarks.
@@ 4.3.2 Maze Navigation Planning
Maze navigation planning [199] involves navigating in a 2D grid world, from the start point to the exit point. Compared with general environment planning, maze planning features a clearly defined start and exit point, as well as a simpler environment. The key challenge lies in multi-hop reasoning while avoiding impassable locations.
@@
Maze navigation problems are often formulated as multiple-choice questions. Depending on the input modality, maze navigation planning can be categorized into three settings: (1) text only, where mazes are represented in ASCII code [212], (2) visual-only, where mazes are depicted as colored images [213, 199, 212], and (3) text-visual, where both ASCII codes and colored images are provided [212].
@@ 4.3.3 Embodied planning
Embodied planning expects agents to leverage multi-modal data to perceive the surrounding environment, generate plans, and take actions in an autonomous manner. General embodied planning tasks are designed around completing a concrete goal. For example, in vision language navigation [214], agents are required to reach a specific location following given instructions, with metrics like success rate and average distance to goal used to measure planning accuracy. By contrast, instead of physically navigating to a goal, embodied planning tasks in VSR require VLMs to solve multi-choice questions, with each option corresponding to possible actions. Depending on the action space, embodied planning in VSR includes vision-and-language navigation and mobile manipulation, both formulated as multi-choice questions.
Vision-and-Language Navigation Vision-and-Language Navigation (VLN) in VSR requires agents to answer multi-choice questions while navigating through the environment. In VSI-Bench [31], the questions are instructions with a cloze (fill-in-the-blank) format, and the answer options correspond to possible actions during the navigation process. In OmniSpatial [33], each question includes the start point, end point, and orientation, while the options represent candidate navigation actions. OmniSpatial [33] also introduces a dynamic analysis setting, where question-answer pairs represent intermediate properties encountered during navigation, such as the number of doors passed.
Mobile Manipulation. In VSR, mobile manipulation tasks ask agents to select the correct option that combines navigation with environment interaction, guided by the provided goals. A representative benchmark for this task is SpaCe-10 [215], which focuses on room tidying tasks and formulates each choice to reflect aspects of the task flow, including navigation paths, goals, goal attributes, and the corresponding actions to be executed.
5 Experiments
TABLE I: Comparison of Benchmarks for Visual Spatial Reasoning.
Benchmark Annotation Scale Input Type QA Type Diversity
(k) Single Multi-View Video T/F Multi-Choice Number
VSR [193] Synthetic 2.81 Concentrated
BLINK [194] Manual 0.28 Diverse
VSI-Bench [31] Synthetic 5.00 Diverse
Spatial-MM [32] Synthetic 2.30 Concentrated
SpatialRGPT [27] Synthetic 1.40 Concentrated
SpatialVLM [26] Synthetic 0.54 Concentrated
STI-Bench [195] Synthetic 2.00 Diverse
SpatialBench [34] Synthetic 0.39 Concentrated
RoboSpatial [216] Synthetic 6.00 Concentrated
What’s up [217] Synthetic 0.49 Concentrated
Space3D-Bench [218] Synthetic 1.00 Concentrated
EmbSpatial-Bench [219] Manual 3.64 Concentrated
OmniSpatial [33] Manual 1.50 Diverse
PhysBench [185] Manual 10.00 Concentrated
VSI-100K [202] Synthetic 5.00 Concentrated
SpatialMQA [220] Manual 5.39 Concentrated
SRBench [210] Synthetic 1.80 Concentrated
Open3DVQA [183] Synthetic 1.24 Concentrated
Q-Spatial Bench [197] Manual 0.27 Concentrated
VoT [213] Synthetic 3.51 Concentrated
3DSRBench [221] Synthetic 2.77 Diverse
Super-CLEVR-3D [190] Synthetic 5.00 Concentrated
SPHERE [186] Manual 2.28 Diverse
LEGO-Puzzles [191] Manual 1.10 Diverse
GEOBench-VLM [222] Manual 10.00 Diverse
OpenEQA [182] Manual 1.60 Concentrated
SPARE3D [206] Synthetic 2.50 Concentrated
VSP [223] Synthetic 4.40 Diverse
DynSuperCLEVR [207] Synthetic 1.00 Concentrated
SITE  [187] Synthetic 8.06 Diverse
SIBench Manual 9.00 Diverse
@@ 5.1 Benchmark Construction
The current benchmarks for VSR exhibit several issues, such as a lack of systematic and comprehensive task design, the presence of low-quality samples, a lack of human-annotated ground truth for some samples, and insufficient task rationale. To address these challenges, we survey existing open-source VSR benchmarks, aiming to integrate high-quality and diverse data. We ultimately develop SIBench, an evaluation benchmark that includes 23 task settings spanning 3 cognitive levels. Table I presents a comparison of current benchmarks with SIBench.
@@
Diversity. To construct a comprehensive evaluation dataset, we investigate more than 20 open-source benchmarks, which we consolidate into 23 high-level task categories (see Fig. 1 for benchmark statistics). These categories span three cognitive levels: basic perception, spatial understanding, and planning. For each task category, we strive to enhance the diversity of the test data. For example, for object size estimation, the data is sourced from three different benchmarks: SPHERE-VLM [186], VSI-Bench [31], and STI-Bench [195]. Our test data includes three types of input formats: single images, multi-view images, and video data. The questions are presented in three formats: multiple-choice, true/false, and video data. Details regarding the SIBench are available in Fig. 1 and Fig. 7.
@@ 5.2 Evaluation
TABLE II: Performance Evaluation of Different Models on SIBench. Bold is the best, and underlining is the second best.
@@
Qwen2.5-VL-7B [41] 10 $0.4172$ 0.5196 0.3946 0.2832
Qwen2.5-VL-72B [41] 4 $0.5114$ 0.5634 0.5019 0.4161
LLaVA-OneVision-7B [18] 7 $0.4844$ 0.5821 0.4644 0.3355
LLaVA-OneVision-72B [18] 5 $0.5103$ 0.6061 0.4889 0.3878
GPT-4o-mini [224] 9 $0.4278$ 0.5505 0.3981 0.3050
Gemini-2.5-Flash [21] 8 $0.4389$ 0.5422 0.3942 0.6100
InternVL-2.5-78B-MPO [41] 6 $0.4983$ 0.5991 0.4635 0.5425
InternVL-3-78B [42] 3 $0.5197$ 0.5947 0.5001 0.4640
InternVL-3.5-38B [225] 2 0.5252 0.5726 0.5134 0.4815
Gemini-2.5-Pro [21] 1 0.5883 0.6425 0.5559 0.8017
TABLE III: Performance Evaluation of Different Models on the SIBench-mini. Bold is the best, and underlining is the second best.
@@
LLaVA-OneVision-72B [18] 7 $0.4987$ 0.5951 0.4633 0.4268
Qwen2.5-VL-72B [41] 6 $0.5006$ 0.6356 0.4526 0.3780
InternVL-3.5-38B [11] 5 $0.5355$ 0.6113 0.5089 0.4878
GLM4.5-V-106B-A12B [22] 4 $0.5822$ 0.6936 0.5404 0.5125
Doubao-Seed-1.6-Vision [226] 3 $0.6216$ 0.6963 0.5922 0.65
Gemini-2.5-Pro [21] 2 0.6295 0.7317 0.5827 0.675
GPT-5 [20] 1 0.6906 0.7248 0.6487 0.775
@@
We perform a detailed comparative analysis of state-of-the-art proprietary and open-source models. Our evaluation of proprietary models includes Gemini-2.5-Pro [21], Gemini-2.5-Flash [21] and GPT-4o-mini [224]. The primary open-source models under consideration are Qwen2.5-VL-72B [41], LLaVA-OneVision-72B [18] and InternVL-3-78B [42]. Consistency is maintained throughout our experimental design. All models process images (both single and multi-view) at an identical resolution. For video inputs, we apply a standard procedure of uniformly sampling 30 frames. Furthermore, we utilize specific prompt templates for each QA category; for example, multiple-choice questions are prefaced with ”Select from the following choices,” and numerical questions are prompted with ”Answer using a number.”
By default, we do not employ Chain-of-Thought (CoT) prompting. To handle any non-formatted outputs, we utilize Gemini-1.5-Flash [21] for post-processing to standardize the answers. The evaluation metric depends on the question type. For True/False (TF) and Multiple-Choice Questions (MCQ), we directly compare the post-processed predictions against the ground truth. For questions requiring a numerical answer, we adopt the Mean Relative Accuracy ($\mathcal{MRA}$) metric from VSI-Bench [31], which is calculated as follows:
$\mathcal{MRA}=\frac{1}{10}\sum_{\theta\in C}\mathbb{I}\left(\frac{|\hat{y}-y|}{y}<1-\theta\right),$ (4)
@@
To streamline the evaluation process, we construct a mini version named SIBench-mini comprising 40 randomly selected samples from each task setting for evaluating Gemini-2.5-Pro [21], Doubao-Seed-1.6-Vision [226], GLM4.5-V-106B-A12B [22], and GPT-5 [20]. This subset is as comprehensive as the full benchmark in terms of task settings, but at a much lower testing cost.
TABLE IV: Performance evaluation of different models on SIBench-mini. Bold is the best, and underlining is the second best.
@@
Height $0.6033$ 0.6400 0.5733 0.7133
Existence 1.0000 0.9500 0.9000 0.9000
Occlusion $0.5600$ 0.6640 0.6160 0.7300
Object Shape 0.3462 0.3077 0.3000 0.4154
Counting 0.5504 0.5084 0.4916 0.4664
Object Size Estimation 0.6085 0.6045 0.6139 0.5968
Spatial Understanding Spatial Compatibility $0.3925$ 0.4019 0.5748 0.7103
Coordinate Conversion 0.6525 0.5950 0.5900 0.6900
Trajectory Description 0.6026 0.5513 0.5513 0.3333
Geometric Reasoning $0.2421$ 0.2579 0.2659 0.3730
Spatial Imagination 0.2950 0.2650 0.3100 0.2700
Spatial Grid 0.8050 0.5200 0.7525 0.9975
Temporal Appearance Order 0.4545 0.3636 0.2727 0.5455
Multi-View Reasoning $0.4125$ 0.4295 0.3998 0.4687
Situational QA 0.5295 0.5095 0.5173 0.5474
Velocity Acceleration 0.4622 0.4403 0.5092 0.4034
Relative Distance $0.5438$ 0.4464 0.5462 0.5459
Camera Pose 0.4306 0.2871 0.5087 0.4041
Spatial Relation 0.8029 0.7884 0.8133 0.6660
Object Localization $0.3769$ 0.7570 0.2461 0.7041
Planning Maze Navigation $0.4350$ 0.3975 0.4850 0.8625
Route Planning $0.2881$ 0.3220 0.4576 0.3898
Overall $0.5114$ 0.5103 0.5252 0.5883
@@
Comparison of Models. The experimental results indicate that leading proprietary models establish the highest tier of performance (see Table II). Specifically, Gemini-2.5-Pro [21] achieves the best performance across the three levels of visual-spatial reasoning included in SIBench. Secondly, while a performance gap persists between open-source models and their leading proprietary counterparts, the former exhibit strong competitiveness against lightweight proprietary models. Although representative open-source models, such as InternVL-3.5-38B [225] (overall score: 0.5252) and Qwen2.5-VL-72B [41] (overall score: 0.5114), lag behind the leading models, their overall performance still outperforms lightweight models like GPT-4o-mini (0.4278) and Gemini-2.5-Flash [21] (0.4389). Furthermore, we observe that Gemini-2.5-Pro holds a significant lead in the dimensions of spatial understanding and planning, which may imply that these premier models possess superior abstract reasoning capabilities. On SIBench-mini, similar patterns are observed. Specifically, as shown in Table III, the proprietary model GPT-5 [20] achieves the best performance, followed by Gemini-2.5-Pro and Doubao-Seed-1.6-Vision [226]. In contrast, other open-source models, such as GLM4.5-V-106B-A12B [22] and InternVL-3.5-38B [11], still exhibit a noticeable performance gap.
@@
Comparison of Various Settings. Our evaluation reveals a stark contrast in the visual spatial capabilities of current models (see Table IV, Table V and Table VI). They demonstrate high proficiency in fundamental perceptual tasks, such as identifying object existence, understanding occlusion, and judging qualitative spatial relationships, indicating strong static feature extraction. However, this proficiency does not extend to higher-level cognitive reasoning. Fig. 1 and Fig. 8 highlight four core deficiencies in the models, which are substantiated by detailed experimental data in the Appendix (see Table V and Table VI). First, the models exhibit poor quantitative reasoning, often failing to accurately estimate physical quantities such as distance and size (see the upper-left part of Fig. 8). This limitation likely arises from their reliance on coarse visual heuristics rather than precise metric representations. Second, they struggle with multi-view reasoning, particularly in tasks requiring the inference of camera pose transformations (see the upper-right part of Fig. 8). This weakness may stem from the lack of explicit geometric constraints in training. Third, their performance degrades significantly on temporal tasks. This fragility in handling non-static inputs suggests their underlying logical inference is unreliable (see the bottom-left part of Fig. 8). A plausible reason is that these models primarily process frames independently, failing to capture coherent temporal dynamics. Finally, and perhaps most critically, the models demonstrate a near-total absence of spatial imagination. All models failed on tasks requiring the creation and manipulation of a mental model (see the bottom-right part of Fig. 8). This exposes a fundamental limitation: they excel at interpreting visible information but cannot conceive of what is absent, functioning as observers rather than reasoners. In short, while models have become adept at ”seeing,” their ability to perform cognitive ”thinking”—particularly spatial imagination and precise inference—remains nascent.

### v2

@@ 4.3.1 Environment Planning
Shortcut Discovery. After observing a video walkthrough from the start to the goal (with unnecessary steps), this task requires VLMs to plan a never-seen shortcut from the start to the goal [SPACE]. The challenges of this task lie in (1) overall understanding of the environment, (2) interactive planning of actions given previous observations.
Route Retracing. In SPACE [SPACE], this task requires VLMs to retrace the route from the start after observing a video from the start to the goal. The route in the video is always the shortest path to the goal. Similar to shortcut discovery, the challenge of this task also lies in the interactive planning of actions.
Map Sketching. After observing a video walkthrough from the start to the goal, map sketching [SPACE] requires VLMs to pick the best map representation of the environment from multiple options. The map contains landmarks, including the start and goal, with different spatial relationships. The challenges of this task lie in (1) recognition of critical landmarks, (2) comprehension of spatial relationships among landmarks.
@@ 4.3.2 Maze Navigation Planning
Maze navigation planning [SPACE] involves navigating in a 2D grid world, from the start point to the exit point. Compared with general environment planning, maze planning features a clearly defined start and exit point, as well as a simpler environment. The key challenge lies in multi-hop reasoning while avoiding impassable locations.
@@
Maze navigation problems are often formulated as multiple-choice questions. Depending on the input modality, maze navigation planning can be categorized into three settings: (1) text only, where mazes are represented in ASCII code [spatialeval], (2) visual-only, where mazes are depicted as colored images [vot, SPACE, spatialeval], and (3) text-visual, where both ASCII codes and colored images are provided [spatialeval].
@@ 4.3.3 Embodied planning
Embodied planning expects agents to leverage multi-modal data to perceive the surrounding environment, generate plans, and take actions in an autonomous manner. General embodied planning tasks are designed around completing a concrete goal. For example, in vision language navigation [wang2024vision], agents are required to reach a specific location following given instructions, with metrics like success rate and average distance to goal used to measure planning accuracy. By contrast, instead of physically navigating to a goal, embodied planning tasks in VSR require VLMs to solve multi-choice questions, with each option corresponding to possible actions. Depending on the action space, embodied planning in VSR includes vision-and-language navigation and mobile manipulation, both formulated as multi-choice questions.
Vision-and-Language Navigation Vision-and-Language Navigation (VLN) in VSR requires agents to answer multi-choice questions while navigating through the environment. In VSI-Bench [VSI-Bench], the questions are instructions with a cloze (fill-in-the-blank) format, and the answer options correspond to possible actions during the navigation process. In OmniSpatial [OmniSpatial], each question includes the start point, end point, and orientation, while the options represent candidate navigation actions. OmniSpatial [OmniSpatial] also introduces a dynamic analysis setting, where question-answer pairs represent intermediate properties encountered during navigation, such as the number of doors passed.
Mobile Manipulation. In VSR, mobile manipulation tasks ask agents to select the correct option that combines navigation with environment interaction, guided by the provided goals. A representative benchmark for this task is SpaCe-10 [gong2025space], which focuses on room tidying tasks and formulates each choice to reflect aspects of the task flow, including navigation paths, goals, goal attributes, and the corresponding actions to be executed.
5 Experiment
@@ 5.1 Benchmark Construction
The current benchmarks for VSR exhibit several issues, such as a lack of systematic and comprehensive task design, the presence of low-quality samples, a lack of human-annotated ground truth for some samples, and insufficient task rationale. To address these challenges, we survey existing open-source VSR benchmarks, aiming to integrate high-quality and diverse data. We ultimately develop SIBench, an evaluation benchmark that includes 23 task settings spanning 3 cognitive levels.
@@
Diversity. To construct a comprehensive evaluation dataset, we investigate nearly 20 open-source benchmarks, which we consolidate into 23 high-level task categories (see Fig. 1 for benchmark statistics). These categories span three cognitive levels: basic perception, spatial understanding, and planning. For each task category, we strive to enhance the diversity of the test data. For example, for object size estimation, the data is sourced from three different benchmarks: SPHERE-VLM [SPHERE], VSI-Bench [VSI-Bench], and STI-Bench [STI-Bench]. Our test data includes three types of input formats: single images, multi-view images, and video data. The questions are presented in three formats: multiple-choice, true/false, and video data. Details regarding the SIBench are available in Fig. 1 and Fig. 7.
@@ 5.2 Evaluation
TABLE I: Performance Evaluation of Different Models on SIBench. Bold is the best, and underlining is the second best.
@@
Qwen2.5-VL-7B [bai2025qwen2] 9 $0.4712$ 0.5280 0.4624 0.3442
Qwen2.5-VL-72B [bai2025qwen2] 4 $0.5447$ 0.5793 0.5430 0.4161
LLaVA-OneVision-7B [li2024llava] 8 $0.4850$ 0.5640 0.4708 0.3355
LLaVA-OneVision-72B [li2024llava] 7 $0.5207$ 0.6086 0.5022 0.3922
GPT-4o-mini [hurst2024gpt] 11 $0.4278$ 0.5505 0.3981 0.3050
Gemini-2.5-Flash [team2023gemini] 10 $0.4389$ 0.5422 0.3942 0.6100
InternVL-2.5-78B-MPO [bai2025qwen2] 5 $0.5338$ 0.5886 0.5165 0.5338
InternVL-3-78B [zhu2025internvl3] 3 $0.5481$ 0.6141 0.5344 0.4488
InternVL-3.5-38B [wang2025internvl3] 6 0.5252 0.5726 0.5134 0.4815
Gemini-2.5-Pro [team2023gemini] 2 0.5883 0.6425 0.5559 0.8017
GPT-5 [gpt5] 1 0.6341 0.6934 0.6158 0.6296
TABLE II: Performance Evaluation of Different Models on the SIBench-mini. Bold is the best, and underlining is the second best.
@@
LLaVA-OneVision-72B [li2024llava] 6 $0.5252$ 0.6136 0.5065 0.3968
Qwen2.5-VL-72B [bai2025qwen2] 7 $0.5168$ 0.5930 0.4958 0.4743
InternVL-3.5-38B [InternVL3] 5 $0.5355$ 0.6113 0.5089 0.4878
GLM4.5-V-106B-A12B [hong2025glm] 4 $0.5822$ 0.6936 0.5404 0.5125
Doubao-Seed-1.6-Vision [guo2025seed1] 3 $0.6216$ 0.6963 0.5922 0.65
Gemini-2.5-Pro [team2023gemini] 2 0.6295 0.7317 0.5827 0.675
GPT-5 [gpt5] 1 0.6906 0.7248 0.6487 0.775
@@
We perform a detailed comparative analysis of state-of-the-art proprietary and open-source models. Our evaluation of proprietary models includes GPT-5 [gpt5], Gemini-2.5-Pro [team2023gemini], Gemini-2.5-Flash [team2023gemini] and GPT-4o-mini [hurst2024gpt]. The primary open-source models under consideration are Qwen2.5-VL-72B [bai2025qwen2], LLaVA-OneVision-72B [li2024llava] and InternVL-3-78B [zhu2025internvl3]. Consistency is maintained throughout our experimental design. All models process images (both single and multi-view) at an identical resolution. For video inputs, we apply a standard procedure of uniformly sampling 30 frames. Furthermore, we utilize specific prompt templates for each QA category; for example, multiple-choice questions are prefaced with ”Select from the following choices,” and numerical questions are prompted with ”Answer using a number.”
By default, we do not employ Chain-of-Thought (CoT) prompting. To handle any non-formatted outputs, we utilize Gemini-1.5-Flash [team2023gemini] for post-processing to standardize the answers. The evaluation metric depends on the question type. For True/False (TF) and Multiple-Choice Questions (MCQ), we directly compare the post-processed predictions against the ground truth. For questions requiring a numerical answer, we adopt the Mean Relative Accuracy ($\mathcal{MRA}$) metric from VSI-Bench [VSI-Bench], which is calculated as follows:
$\mathcal{MRA}=\frac{1}{10}\sum_{\theta\in C}\mathbb{1}\left(\frac{|\hat{y}-y|}{y}<1-\theta\right),$ (4)
@@
To streamline the evaluation process, we construct a mini version named SIBench-mini comprising 40 randomly selected samples from each task setting for evaluating Gemini-2.5-Pro [team2023gemini], Doubao-Seed-1.6-Vision [guo2025seed1], GLM4.5-V-106B-A12B [zeng2025glm], and GPT-5 [gpt5]. This subset is as comprehensive as the full benchmark in terms of task settings, but at a much lower testing cost.
TABLE III: Performance evaluation of different models on visual spatial reasoning tasks. Bold is the best, and underlining is the second best.
@@
Height $0.6000$ 0.6466 0.6800 0.7133
Existence 1.0000 0.9500 0.9250 0.9000
Occlusion $0.5580$ 0.6660 0.6460 0.7300
Object Shape $0.3538$ 0.3077 0.5154 0.4154
Counting 0.5504 0.5084 0.7408 0.4664
Object Size Estimation 0.6085 0.6045 0.7645 0.5968
Spatial Understanding Spatial Compatibility 0.6028 0.5841 0.5607 0.7103
Coordinate Conversion $0.6513$ 0.5950 0.7788 0.6900
Trajectory Description 0.6026 0.5513 0.3205 0.3333
Geometric Reasoning $0.2421$ 0.2579 0.5278 0.3730
Spatial Imagination 0.2900 0.2650 0.4575 0.2700
Spatial Grid 0.7875 0.5175 0.7900 0.9975
Temporal Appearance Order 0.5455 0.3636 0.2957 0.5455
Multi-View Reasoning $0.4072$ 0.4316 0.5854 0.4687
Situational QA $0.5295$ 0.5128 0.6055 0.5474
Velocity Acceleration 0.4605 0.4403 0.5746 0.4034
Relative Distance 0.5675 0.5037 0.6828 0.5459
Camera Pose 0.4738 0.3169 0.5000 0.4041
Spatial Relation 0.7978 0.7870 0.6929 0.6660
Object Localization $0.7134$ 0.7570 0.7727 0.7041
Planning Maze Navigation $0.4350$ 0.3975 0.7725 0.8625
Route Planning $0.2881$ 0.3559 0.3351 0.3898
Overall $0.5447$ 0.5207 0.6341 0.5883
@@
Comparison of Models. The experimental results indicate that leading proprietary models establish the highest tier of performance (see Table I). Specifically, GPT-5 [gpt5] achieves the best performance across the three levels of visual-spatial reasoning included in SIBench and SIBench-mini. Secondly, while a performance gap persists between open-source models and their leading proprietary counterparts, the former exhibit strong competitiveness against lightweight proprietary models. Although representative open-source models, such as InternVL-3.5-38B [wang2025internvl3] (overall score: 0.5252) and Qwen2.5-VL-72B [bai2025qwen2] (overall score: 0.5114), lag behind the leading models, their overall performance still outperforms lightweight models like GPT-4o-mini (0.4278) and Gemini-2.5-Flash [team2023gemini] (0.4389). Furthermore, we observe that GPT-5 and Gemini-2.5-Pro holds a significant lead in the dimensions of spatial understanding and planning, which may imply that these premier models possess superior abstract reasoning capabilities. On SIBench-mini, similar patterns are observed. Specifically, as shown in Table II, the proprietary model GPT-5 [gpt5] achieves the best performance, followed by Gemini-2.5-Pro and Doubao-Seed-1.6-Vision [guo2025seed1]. In contrast, other open-source models, such as GLM4.5-V-106B-A12B [zeng2025glm] and InternVL-3.5-38B [InternVL3], still exhibit a noticeable performance gap.
@@
Comparison of Various Settings. Our evaluation reveals a stark contrast in the visual spatial capabilities of current models (see Table III, Table IV and Table V). They exhibit high proficiency in fundamental perceptual tasks, such as identifying object existence, understanding occlusion, and judging qualitative spatial relationships, indicating strong static feature extraction. However, this proficiency does not extend to higher-level cognitive reasoning. Fig. 1 and Fig. 8 highlight four core deficiencies in the models, which are substantiated by detailed experimental data in the Appendix (see Table IV and Table V). First, the models exhibit poor quantitative reasoning, often failing to accurately estimate physical quantities such as distance and size (see the upper-left part of Fig. 8). This limitation likely arises from their reliance on coarse visual heuristics rather than precise metric representations. Second, they struggle with multi-view reasoning, particularly in tasks requiring the inference of camera pose transformations (see the upper-right part of Fig. 8). This weakness may stem from the lack of explicit geometric constraints in training. Third, their performance degrades significantly on temporal tasks. This fragility in handling non-static inputs suggests their underlying logical inference is unreliable (see the bottom-left part of Fig. 8). A plausible reason is that these models primarily process frames independently, failing to capture coherent temporal dynamics. Finally, and perhaps most critically, the models exhibit a near-total absence of spatial imagination. All models failed on tasks requiring the creation and manipulation of a mental model (see the bottom-right part of Fig. 8). This exposes a fundamental limitation: they excel at interpreting visible information but cannot conceive of what is absent, functioning as observers rather than reasoners. In short, while models have become adept at ”seeing,” their ability to perform cognitive ”thinking”, particularly spatial imagination and precise inference, remains nascent.